OpenStudy (jack_mrcus):

What is the equation in standard form of the line which passes through (1, -3) and has a slope of 2?

OpenStudy (karim728):

first find the slope(m) using (y2-y1)/(x2-x1) then use this y-y1=m(x-x1)

OpenStudy (jack_mrcus):

so solve for y-(-3)=2(x-1) ?

OpenStudy (jack_mrcus):

so y=x-5?

OpenStudy (karim728):